# {textblock*}

### v2

@@ V Conclusions
This work systematically analyzed the importance and design of the various building blocks that compose speech-to-text models by connecting an SFM encoder and an LLM through an adapter. To this aim, we compared all the combinations of 2 SFMs, 2 LLMs, and 5 adapters, which mostly differ for their length reduction module. With comprehensive experiments covering two tasks – ASR and ST – and 5 language directions, our results demonstrate that the choice of the SFM is the most critical factor influencing downstream performance. We also established that there is no one-size-fits-all solution for the length adapter, as the optimal choice varies depending on the specific combination of SFM and LLM. Notably, the Base and WLQ-former adapters, which feature very different compression factors, demonstrate strong performance across tasks, suggesting that reducing sequence length mismatch between speech and text is less crucial than previously assumed.
Acknowledgments
This paper has received funding from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People). We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid ACC Cyfronet AGH for providing computer facilities and support within computational grant no PLG/2024/016971. Marco Gaido was supported by the PNRR project FAIR - Future AI Research (PE00000013), under the NRRP MUR program funded by the NextGenerationEU. This work has been carried out while Pierfrancesco Melucci was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with Università di Roma Tre.

### v3

@@ 5 Conclusions
This work systematically analyzed the importance and design of the various building blocks that compose speech-to-text models by connecting an SFM encoder and an LLM through an adapter. To this aim, we compared all the combinations of 2 SFMs, 2 LLMs, and 5 adapters, which mostly differ in their length reduction module. With comprehensive experiments covering two tasks–ASR and ST–and 5 language directions, our results demonstrate that the choice of the SFM is the most critical factor influencing downstream performance. We also established that there is no one-size-fits-all solution for the length adapter, as the optimal choice varies depending on the specific combination of SFM and LLM. Notably, the Base and WLQ-former adapters, which feature very different compression factors, demonstrate strong performance across tasks, suggesting that reducing sequence length mismatch between speech and text is less crucial than previously assumed.
6 Acknowledgments
This paper has received funding from the European Union’s Horizon research and innovation programme under grant agreement No 101135798, project Meetween (My Personal AI Mediator for Virtual MEETtings BetWEEN People).
We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid ACC Cyfronet AGH for providing computer facilities and support within computational grant no PLG/2024/016971.
Marco Gaido was supported by the PNRR project FAIR - Future AI Research (PE00000013), under the NRRP MUR program funded by the NextGenerationEU.
This work has been carried out while Pierfrancesco Melucci was enrolled in the Italian National Doctorate on Artificial Intelligence run by Sapienza University of Rome in collaboration with Università di Roma Tre.